OpenStudy (anonymous):

M(8, 7) is the midpoint of RS. The coordinates of S are (9, 5). What are the coordinates of R?

OpenStudy (anonymous):

Can anyone please tell me how to find the Coordinates of R?

OpenStudy (leader):

9+x/2=8 9+x=16 x=7 y+5/2=7 5+y=14 y=9 So the point is (7,9)

OpenStudy (leader):

@vic21

OpenStudy (anonymous):

Thank you :)